# Hardware test and validation of the angular droop control: Analysis and experiments

### v1

@@ V-D Experimental Results
    Figure 16: Duty cycles of the indirect (a) and direct control (b) schemes in the black start experiment.
VI Scenario II: Two identical converters to a common load
VI-A Scenario description
This scenario consists of two identical three-phase balanced DC/AC converters supplied by two independent DC sources behind DC/DC converters and connected to a common resistive load as shown in Fig. 17. The control scheme of the boost converters follow the vector control discussed in Section III-B1. The DC/AC converters are in closed-loop with the angular droop control according to the direct implementation described in Section III-B2. Table II summarizes the parameter values. To satisfy the modeling assumptions in Section II-A, we set the ratio $X/R_{\ell}\approx 11$ with the reactance $X=\omega^{*}L_{\ell}$, corresponding to highly inductive power network [34, 26].
TABLE II: Parameters values of the transmission line.
Symbol Definition Range Value in S.I.
$R_{l}$ line resistance >0 $20\cdot 10^{-3}$
$L_{l}$ line inductance >0 $700\cdot 10^{-6}$
Figure 17: Schematic representation of Scenario II consisting in two identical converter systems, each in closed-loop with angular droop control, connected to a common resistive load via highly inductive transmission lines. The line resistance is negligible.
VI-B Experimental procedure
We start by connecting the DC/AC Converter I following the steps described in Section V-B) to the resistive load. From this initial setup, we conduct the following experiments:
  1. 1.
    Frequency synchronization: we connect the DC/AC Converter II at $t=0$ to the initial setup consisting in Converter I connected to resistive load by closing the connection relay and enabling its modulation.
  2. 2.
    Power sharing: Starting from a frequency synchronous two-converter system, we study in experiments the power sharing capabilities of the angular droop control by suitable tuning of the power-to-angle droop gains.
VI-C Challenge 2: Clock drift in angular droop controllers
VI-C1 Analysis
Angular droop control is susceptible to clock drifts. This is well-documented in the power system literature [26, 35, 36, 37] and can be explained as follows. Let $t_{k}>0$ denote the local time at the $k$-th DC/AC converter with respect to a global reference time $t$ as follows,
@@ VI-C1 Analysis
$P^{s}_{k}$ $=\frac{V_{k}\,V_{j}}{X_{kj}}\sin(\theta^{s}_{k}-\theta^{s}_{j})=\frac{V_{k}\,V_{j}}{X_{kj}}\sin((\omega_{k}-\omega_{j})t)$
$=\frac{V_{k}\,V_{j}}{X_{kj}}\sin(\omega^{*}\,t\,(\epsilon_{k}-\epsilon_{j})),$
where $\theta^{s}_{j}(t):=\omega_{j}\,t$ is the steady state angle, $\omega_{j}$ is the local frequency at the $j-$th converter and $X_{jk}$ is the reactance between converters $k$ and $j$ following the Kron-reduction. It can be deduced that when local clock drifts are not compensated for, the injected active power $P^{s}_{k}$ drifts apart from its nominal value. Since the angular droop control law involves an integration, the closed-loop dynamics (3) are not robust to local clock drifts [37] and suitable solution needs to be developed for the hardware implementation.
VI-C2 Proposed solution
In our lab experiments, we utilize Imperix distributed modulation allowing for synchronized PWM signals generation across multiple digital controllers [39]. This allows multiple controllers to behave as if all PWM signals were generated from a single centralized modulator. The distributed modulators must provide the same time resolution and accuracy as modulators belonging to a unique hardware. This technology consists in the distribution of a common high-frequency clock (or master clock) across the entire control network through a direct optical fiber connection between the clocks of Converter I and II [40]. Therefore, the distributed devices belong to same clock domain, which eliminates the clock drift for the integration actions and allows for distributed modulation with an accuracy of ±$\pm 2\,\mathrm{ns}$.
VI-D Experimental results
Figure 18: Frequency (a) and angle error (b) for $k=1,2$ and angle differences (c), following the connection of Converter II to Converter I at $t=0$, both in closed-loop with the direct angular droop control for $\gamma_{k}=5\cdot 10^{4}$ and $\alpha_{k}=2000$ for $k=1,2$. Here $\theta^{*}_{1}(t)=\omega^{*}t$ for $t<0$ and $\theta^{*}_{2}(t)=\omega^{*}t+\theta_{1}(0)-0.042$ for $t\geq 0$, where $\theta_{1}(0)$ is the modulation angle of Converter I at the time of interconnection.
@@ VI-D Experimental results
VI-D1 Frequency synchronization
Figure 18 depicts the frequency and angle errors following the connection of Converter I to Converter II. Before the interconnection, the modulation angle of Converter I is initialized at zero with $\theta_{1}^{*}(t)=\omega^{*}t$, for $t<0$. The nominal angle of Converter II is given by $\theta^{*}_{2}(t)=\omega^{*}t+\theta_{2}(0),\,t\geq 0$, where $\theta_{2}(0)$ is the initial angle of Converter II at the time of interconnection $t=0$. The choice of the initial angle $\theta^{*}_{2}(0)$ can be determined as follows. The two-converter system represented in Fig. 17 can be reduced to two sources connected to one load as depicted in Fig. 19. The active power at the $k$-th DC/AC converter at steady state is given by [2]
@@ VI-D1 Frequency synchronization
where $V_{0}\angle\theta^{s}_{0}$ is the phasor at the common node connecting the two converters and $V_{k}\angle\theta^{s}_{k}$ that of the switching voltage of the $k-$th DC/AC converter with $V_{k}=\lVert\frac{1}{2}u_{k}V_{dc}\rVert$ and $X_{k0}=\omega^{*}L_{l}>~{}0$ is the reactance of the inductive line impedance. Thus, the active power at Converter II is given by $P^{s}_{2}=\frac{V_{2}\,V_{0}}{X_{20}}\sin{(\theta^{s}_{2}-\theta^{s}_{0})}$. Setting $P^{s}_{2}=P^{*}_{2}=0$ leads to $\theta_{2}(0)=\theta_{0}(0)$. From $P^{s}_{1}=\frac{V_{1}\,V_{0}}{X_{10}}\sin{(\theta^{s}_{1}-\theta^{s}_{0})}=P^{*}$, we obtain
$\theta_{1}(0)=\theta_{0}(0)+\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}}),$ (26)
where $\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}})=0.006$, $\theta_{1}(0)$ is the initial angle of Converter I and $\theta_{0}(0)$ is the initial angle at node 0 depicted in Fig. 19 obtained from a Phased-Locked-Loop (PLL) scheme [41] at the time of interconnection. Our experimental results in Fig. 18 show that the modulation angle differences at steady state, i.e., $\theta^{s}_{1}-\theta^{s}_{2}=0.002$. This can be inferred from the negligible but non-zero line resistance $R_{l}>0$ leading to inevitable power mismatches, also visible in Fig. 20, as follows
$\theta_{1}^{s}-\theta_{2}^{s}$ $=\theta_{1}^{*}-\theta_{2}^{*}+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$
$=\theta_{1}(0)-\theta_{0}(0)+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$
$=\theta_{0}(0)+\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}})-\theta_{0}(0)+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$
$=\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}})+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$
$=0.006+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2}),$
where $\delta P_{1}+\delta P_{2}=-200$ and $\delta P_{1}=P^{*}_{1}-P^{s}_{1}$ and $\delta P_{2}=P^{*}_{2}-{P}^{s}_{2}$. Thereby, the steady state angle differences remain within $[-\pi/2,\pi/2]$ (rad) and the security constraint in Assumption 1 is satisfied. Under Assumption 1, the two converters synchronize at nominal frequency $\omega^{*}$ within $0.25\,s$. Both converters’ angles converge to their frequency synchronous steady states. This corresponds to the active power of Converter I and II, $P^{s}_{1},P^{s}_{2}$ reaching a nearby nominal value as seen in Figure 20. We note thereby that $P^{*}_{1}+P^{*}_{2}=P^{*}$ where $P^{*}$ is the total active power drawn by the load resistance. Furthermore, the converters exchange reactive power $\tilde{Q}^{s}_{1}<0$ and $\tilde{Q}^{s}_{2}>0$, where $\tilde{Q}^{s}_{1}+\tilde{Q}^{s}_{2}=0$ at all times $t\geq 0$. This can be explained by the unavoidable mismatches of the voltage amplitudes (see also Fig. 21), which leads to non-zero reactive power. In particular, in a highly inductive, Kron-reduced [38] power network, the reactive power is expressed at the $k-$th converter by [42]
@@
Hence, $\tilde{Q}^{s}_{k}\neq 0$ for $k=\{1,2\}$ due to $V_{k}\neq V_{j}$. Finally, the phase portrait of the output voltages $v_{1}$ and $v_{2}$ in Fig. 21 represents a limit cycle of an approximate radius of $V^{d}$ in the phase plane which shows once again the frequency synchronization in AC voltages nearby a desired voltage amplitude $V^{d}$. As a conclusion, our experiment validates the local asymptotic stability result shown in [20].
@@
Figure 21: Phase portrait of the periodic orbit of the three-phase output voltages $v_{1}$ and $v_{2}$ in $abc$- frame, respectively in closed-loop with the direct angular droop control for $\gamma_{k}=5\cdot 10^{4}$ and $\alpha_{k}=2000$ and $k=1,2$ during the frequency synchronization experiment.
VI-D2 Power sharing
Figure 22: Frequency (a), angle error (b) for $k=1,2$ and angle differences (c), following the connection of Converter II to Converter I at $t=0$, both in closed-loop with the direct angular droop control for $\gamma_{k}=500$ and $\alpha_{k}=2000$ for $k=1,2$ and $r=1$ in the power sharing experiment.
Figure 23: Active power $P_{k}=v_{k}^{\top}i_{o,k}$ in (a) and reactive power $Q_{k}=v_{k}^{\top}\mathbf{J}i_{o,k}$ in (b) following the connection of Converter II to Converter I at $t=0$, both in closed-loop with the direct angular droop control for $\gamma_{k}=500,\alpha_{k}=2000$ and the power ratio $r=1$ in (33) for $k=1,2$ in the power sharing experiment, $P_{1}^{*}=P^{*}_{2}=\frac{P^{*}}{2}$.
To achieve power sharing among the Converters I and II, we first determine a suitable tuning for the power-to-angle droop gains $\gamma_{k},\,k=1,2$ and the nominal power ratio $P_{1}^{*}/P_{2}^{*}$, by conducting the following analysis inspired by [26]. To keep the analysis tractable, we assume in the remainder an inductive power network characterized by high ${X}/{R_{l}}$ ratio (see Table II), constant voltage magnitudes and neglect the output filter at each converter.
@@ VI-D2 Power sharing
$\theta^{*}_{1}$ $-\theta^{*}_{2}+\frac{1}{\gamma_{1}}P_{1}^{*}-\frac{1}{\gamma_{2}}P_{2}^{*}$
$=\bigg{(}\frac{1}{\gamma_{1}}+\frac{X_{10}}{V_{1}\,V_{0}}\bigg{)}P^{s}_{1}-\bigg{(}\frac{1}{\gamma_{2}}+\frac{X_{20}}{V_{2}\,V_{0}}\bigg{)}P^{s}_{2}.$
@@
holds, it yields that
$\frac{1}{\gamma_{1}}P_{1}^{*}-\frac{1}{\gamma_{2}}P_{2}^{*}\approx\frac{1}{\gamma_{1}}P^{s}_{1}-\frac{1}{\gamma_{2}}P^{s}_{2},$
and we deduce that for an active power ratio defined as,
$r:=\frac{P^{*}_{1}}{P^{*}_{2}}=\frac{\gamma_{1}}{\gamma_{2}},$ (33)
it holds that,
$\frac{P^{s}_{1}}{P^{s}_{2}}\approx r,$ (34)
and the power sharing amongst the two converters is guaranteed. Finally, for our experimental setup, the condition (32) can be rewritten as
@@
with $V_{1}=V_{2}=V^{d},V_{0}\approx V^{d}$ and $X_{10}=X_{20}=\omega^{*}L_{l}$. To achieve (35), we select $\gamma_{1}=\gamma_{2}=500$ with $r=1$ throughout the power sharing experiment.

### v2

@@ III-C3 Comparison between direct and indirect control
Figure 14: Duty cycles $d_{AC}$ of the indirect (a) and direct control (b) schemes in the black start experiment.
TABLE II: Comparison between direct and indirect control. RMS denotes the root mean square, nadir the smallest value, max. the maximum value and $T_{\mathrm{set}}$ the settling time to $\pm 0.02\mathrm{Hz}$ frequency deviation following the event.
Black start experiment
control nadir $\Delta\omega$ RMS $\Delta\omega$ max. $|\Delta\theta|$ RMS $\Delta\theta$ $T_{\mathrm{set}}$ direct -0.159 0.165 0.04 0.024 0.61 indirect -0.173 0.181 0.48 0.027 0.61
Load step experiment
control nadir $\Delta\omega$ RMS $\Delta\omega$ max. $|\Delta\theta|$ RMS $\Delta\theta$ $T_{\mathrm{set}}$ direct -0.485 0.082 0.018 0.0178 0.11 indirect -0.162 0.045 0.019 0.0170 0.24
IV Scenario II: Two identical converters to a common load
The use of angles instead of frequency has further implications for the real-world application of angular droop in grids with multiple converters. Unlike nominal frequency, which is constant, the nominal angle changes over time. This must be considered for the parallel operation of distributed DC/AC converters controlled by angular droop.
IV-A Scenario description
Scenario II consists of two identical DC/AC converter systems supplied by two independent DC sources behind DC/DC converters and connected to a common resistive load as shown in Fig. 15. Table I summarizes the parameter values of the transmission line replicas. To satisfy the modeling assumption of highly inductive transmission lines, we set the ratio $X/R_{\ell}\approx 11$ with the reactance $X=\omega^{*}L_{\ell}$, see also [23]. In this scenario, we verify:
  • frequency synchronization capabilities of the angular droop control.
  • power-sharing capabilities in dependence of the angular droop control gains.
We hereby underscore the relevance of the two-converter test case as a toy example that provides a solid foundation for generalizations towards $n$-converter system with $n>2$ for the following reasons. First, frequency synchronization and power-sharing are fundamental properties that are shared by a network of converter of any size $n>1$. Second, any analysis comprising $n$-converters with $n>2$ can be divided into a pairwise study of two converter $i$ and $j$ and therefore reduces to a two-converter setup. Third, the same implementation challenges, in particular, clock drifts are encountered in any extension to a network.
Figure 15: Schematic representation of Scenario II consisting in two identical converter systems, each in closed-loop with angular droop control, connected to a common resistive load via highly inductive transmission lines.
IV-B Challenge: Clock drift in angular droop controlled grids
Analysis
Angular droop control is susceptible to clock drifts. This is well-documented in the power system literature [23, 29] and can be explained as follows. Let $t_{k}>0$ denote the local time at the $k$-th DC/AC converter with respect to a global reference time $t$ as follows,
@@ IV-B Challenge: Clock drift in angular droop controlled grids
$P^{s}_{k}$ $=\frac{V_{k}\,V_{j}}{X_{kj}}\sin(\theta^{s}_{k}-\theta^{s}_{j})=\frac{V_{k}\,V_{j}}{X_{kj}}\sin((\omega_{k}-\omega_{j})t)$ (19)
$=\frac{V_{k}\,V_{j}}{X_{kj}}\sin(\omega^{*}\,t\,(\epsilon_{k}-\epsilon_{j})),$ (20)
where $\theta^{s}_{j}(t):=\omega_{j}\,t$ is the steady state angle, $\omega_{j}$ is the local frequency at the $j$-th converter and $X_{jk}$ is the reactance between converters $k$ and $j$ following Kron-reduction. It can be deduced that when local clock drifts are not compensated for, the injected active power $P^{s}_{k}$ drifts apart from its nominal value. Since the angular droop control law involves an integration, the closed-loop dynamics (3) are not robust to local clock drifts [29] and suitable solution needs to be developed for the hardware implementation.
Proposed solution
In our lab experiments, we utilize the distribution of a common high-frequency clock [27] (or master clock) across the entire control network through a direct optical fiber connection between the controllers of Converter I and II [13]. Therefore, the distributed devices belong to same clock domain, which eliminates the clock drift for the integration actions. Our proposed solution enables a simple yet resource-friendly realization of angular droop control for testing multi-converter cases and provides a synchronization accuracy of $\pm 2\mathrm{ns}$ [27], surpassing the clock accuracy of the global positioning system (GPS), which is in the range of $\pm 10\,\mathrm{ns}$ to $\pm 100\,\mathrm{ns}$ [13].
IV-C Experimental results
IV-C1 Frequency synchronization
Figure 16 depicts the frequency and angle errors following the connection of Converter I and Converter II. Before the interconnection, the modulation angle of Converter I is initialized at zero with $\theta_{1}^{*}(t)=\omega^{*}t$, for $t<0$ and $\theta_{1}^{*}(t)=\omega^{*}t+\theta_{1}^{*}(0)$, for $t\geq 0$. The nominal angle of Converter II is given by $\theta^{*}_{2}(t)=\omega^{*}t+\theta^{*}_{2}(0),\,t\geq 0$, where $\theta^{*}_{1}(0)$ and $\theta^{*}_{2}(0)$ are the initial angle of Converter I and II at the time of interconnection $t=0$. The choice of the initial angle $\theta^{*}_{2}(0)$ can be determined as follows. The two-converter system represented in Fig. 15 can be reduced to two sources connected to one load as depicted in Fig. 17.
Figure 16: Frequency (a) and angle error (b) for $k=1,2$ and angle differences (c), following the connection of Converter II to Converter I at $t=0$, both in closed-loop with the direct angular droop control for $\gamma_{k}=5\cdot 10^{4}$ and $\alpha_{k}=2000$ for $k=1,2$. Here $\theta^{*}_{1}(t)=\omega^{*}t$ for $t<0$ and $\theta^{*}_{2}(t)=\omega^{*}t+\theta_{2}^{*}(0)$ for $t\geq 0$, where $\theta_{1}(0)$ is the modulation angle of Converter I at the time of interconnection.
@@ IV-C1 Frequency synchronization
The active power at the $k$-th DC/AC converter at steady state is given by [24]
@@
where $V_{0}\angle\theta^{s}_{0}$ is the phasor at the common node connecting the two converters and $V_{k}\angle\theta^{s}_{k}$ that of the switching voltage of the $k$-th DC/AC converter with $V_{k}=\lVert\frac{1}{2}u_{k}V_{dc}\rVert$ and $X_{k0}=\omega^{*}L_{l}>~0$ is the reactance of the inductive line impedance. Thus, the active power at Converter II is given by $P^{s}_{2}=\frac{V_{2}\,V_{0}}{X_{20}}\sin{(\theta^{s}_{2}-\theta^{s}_{0})}$. Setting $P^{s}_{2}=P^{*}_{2}=0$ leads to $\theta^{*}_{2}(0)=\theta_{0}(0)$. From $P^{s}_{1}=\frac{V_{1}\,V_{0}}{X_{10}}\sin{(\theta^{s}_{1}-\theta^{s}_{0})}=P^{*}$, we obtain
$\theta^{*}_{1}(0)=\theta_{0}(0)+\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}}),$ (22)
where $\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}})=0.006$, $\theta_{1}(0)$ is the initial angle of Converter I and $\theta_{0}(0)$ is the initial angle at node 0 depicted in Fig. 17 obtained from a Phased-Locked-Loop (PLL) scheme at the time of interconnection.
Our experimental results in Fig. 16 show that the modulation angle differences at steady state, i.e., $\theta^{s}_{1}-\theta^{s}_{2}=0.002$. This can be inferred as follows
$\theta_{1}^{s}-\theta_{2}^{s}$ $=\theta_{1}^{*}-\theta_{2}^{*}+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$ (23)
$=\theta_{1}(0)-\theta_{0}(0)+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$ (24)
$=\theta_{0}(0)+\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}})-\theta_{0}(0)+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$ (25)
$=0.006+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2}),$ (26)
where $\delta P_{1}+\delta P_{2}=-200$ and $\delta P_{1}=P^{*}_{1}-P^{s}_{1}$ and $\delta P_{2}=P^{*}_{2}-{P}^{s}_{2}$. Thereby, the steady state angle differences remain within $\pi/2,\pi/2]$ (rad) and the security constraint in [17] is satisfied and the two converters synchronize at nominal frequency $\omega^{*}$ within 0.25s0.25\,s. Both converters’ angles converge to their frequency synchronous steady states. This corresponds to the active power of Converter I and II, $P^{s}_{1},P^{s}_{2}$ reaching a nearby nominal value as seen in Figure 18. We note thereby that $P^{*}_{1}+P^{*}_{2}=P^{*}$ where $P^{*}$ is the total active power drawn by the load resistance. Furthermore, the converters exchange reactive power $\tilde{Q}^{s}_{1}<0$ and $\tilde{Q}^{s}_{2}>0$, where $\tilde{Q}^{s}_{1}+\tilde{Q}^{s}_{2}=0$ at all times $t\geq 0$. In particular, in a highly inductive, Kron-reduced [8] power network, the reactive power is expressed at the $k$-th converter by [24]
@@
Hence, mismatches of the voltage amplitudes, $V_{k}\neq V_{j}$ (see also Fig. 19) lead to non-zero reactive power $\tilde{Q}^{s}_{k}\neq 0$ for $k=\{1,2\}$. Such mismatches are unavoidable in hardware setups due to manufacturing tolerances of the used components and parasitic effects that occur in real-world applications. Finally, the phase portrait of the output voltages $v_{1}$ and $v_{2}$ in Fig. 19 represents a limit cycle of an approximate radius of $V^{d}$ in the phase plane which shows once again the frequency synchronization in AC voltages nearby a desired voltage amplitude $V^{d}$. As a conclusion, our experiment validates local asymptotic stability result shown in [17].
@@
Figure 19: Phase portrait of the periodic orbit of the three-phase output voltages $v_{1}$ and $v_{2}$ in abcabc- frame, respectively in closed-loop with the direct angular droop control for $\gamma_{k}=5\cdot 10^{4}$ and $\alpha_{k}=2000$ and $k=1,2$ during the frequency synchronization experiment.
IV-C2 Power-sharing
Figure 20: Active power $P_{k}=v_{k}^{\top}i_{o,k}$ in (a) and reactive power $Q_{k}=v_{k}^{\top}\mathbf{J}i_{o,k}$ in (b) following the connection of Converter II to Converter I at $t=0$, both in closed-loop with the direct angular droop control for $\gamma_{k}=500,\alpha_{k}=2000$ and the power ratio $r=1$ in (34) for $k=1,2$ in the power-sharing experiment, $P_{1}^{*}=P^{*}_{2}=\frac{P^{*}}{2}$.
To achieve power-sharing among the Converters I and II, we first determine a suitable tuning for the power-to-angle droop gains $\gamma_{k},\,k=1,2$ and the nominal power ratio $P_{1}^{*}/P_{2}^{*}$, by conducting the following analysis inspired by [23]. To keep the analysis tractable, we assume in the remainder an inductive power network characterized by high $X/{R_{l}}$ ratio (see Table I), constant voltage magnitudes and neglect the output filter at each converter.
@@ IV-C2 Power-sharing
$\theta^{*}_{1}$ $-\theta^{*}_{2}+\frac{1}{\gamma_{1}}P_{1}^{*}-\frac{1}{\gamma_{2}}P_{2}^{*}=\bigg(\frac{1}{\gamma_{1}}+\frac{X_{10}}{V_{1}\,V_{0}}\bigg)P^{s}_{1}-\bigg(\frac{1}{\gamma_{2}}+\frac{X_{20}}{V_{2}\,V_{0}}\bigg)P^{s}_{2}.$ (32)
@@
holds, it yields that $\frac{1}{\gamma_{1}}P_{1}^{*}-\frac{1}{\gamma_{2}}P_{2}^{*}\approx\frac{1}{\gamma_{1}}P^{s}_{1}-\frac{1}{\gamma_{2}}P^{s}_{2}$, and we deduce that for an active power ratio defined by,
$r:=\frac{P^{*}_{1}}{P^{*}_{2}}=\frac{\gamma_{1}}{\gamma_{2}}.$ (34)
Therefore, $\frac{P^{s}_{1}}{P^{s}_{2}}\approx r$ and the power-sharing between the two converters is guaranteed. Finally, for our experimental setup, the condition (33) can be rewritten as
@@
with $V_{1}=V_{2}=V^{d},V_{0}\approx V^{d}$ and $X_{10}=X_{20}=\omega^{*}L_{l}$. To achieve (35), we select $\gamma_{1}=\gamma_{2}=500$ with $r=1$ throughout the power-sharing experiment.